# Concentration Inequalities for Branching Random Walks

### v5

@@ 1.1 Concentration inequalities
For a long time, there had been no substantive improvement in the eliminating boundedness condition (the trivial idea of truncation gives weak results, though) [^2] In fact, it was not until 1995 that the first improvement in Hoeffding inequality was made [42]. The reason might be two folds. First, there lacked motivation. Second, nifty techniques used in the pioneer work and the later literatures as well to achieve finer results somehow concealed the essential reason why the inequalities hold.. In early 2000s [39] relaxed the boundedness to finite exponential moments; i.e. for a $\delta>0,\mathbb{E}(e^{\delta X_{i}})$ is uniformly bounded for all $i$, and gave $\Pr(|S_{n}-\mathbb{E}(S_{n})|\geq\lambda|)\leq e^{-c\big(\frac{\lambda^{2}}{n}\big)^{1/3}}$— not Chernoff bound. Later, motivated by the study of the free energy of a directed polymer in a random environment, [34] gave Chernoff bound under the hypothesis that, for a $\delta>0,\mathbb{E}(e^{\delta X_{i}}\mid\mathcal{F}_{i-1})$ is bounded for all $i$ (where $\mathcal{F}_{i-1}$ is a filtration for the martingale difference sequence $X_{i}$). Other improvements of concentration inequalities in relaxing boundedness condition can be found in [8] [20] [26] [21]. However, to our knowledge, no existing work eliminates both the martingale and boundedness conditions while achieving sub-Gaussian concentration; related work by Wormald (1995) is discussed later. The first major contribution of this paper is to do so. Informally, the small per-step effect in a random process can be controlled by the Lipschitz continuity of the mean increment with respect to the history, not necessarily by “almost independence”.
@@ 1.2 Notation and contributions
$\frac{\partial y}{\partial t}:=\frac{\mathbb{E}_{i}(y_{t})-\mathbb{E}_{i}(y_{t-1})}{1/N}=\mathbb{E}_{i}(Y_{t})-\mathbb{E}_{i}(Y_{t-1})=E_{i}(X_{t})$ (1.2)
@@
While, traditionally, concentration inequalities require the condition (1.4), one of the primary contributions of this work is the relaxation of this classical constraint to the condition:
@@ 2 Extension of Hoeffding-Azuma inequality
$\exp({\sim nt^{2}-t\lambda})$
which, given $\lambda=o(n)$, is exponentially small, $\exp(-n^{p})$ with $0<p<1$, under the choice $t\sim n^{(p-1)/2}$.
@@
The emphasis in this exposition is to show that the condition of $X_{i}<\infty$ can be relaxed to $E(e^{|X_{i}|})<\infty$ (owing to [34] and [43]) without compromising strength of results; from the practical point of view, the difference between the constant factors in the exponent, i.e., $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}$ versus $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$, is immaterial – both are square-exponential decay. In addition, it is shown that sub-Gaussian tail bounds can be obtained without “nifty” techniques.
Next, we show what weaker conditions can ensure concentration inequalities, beyond independence, martingale differences, and boundedness. To this end, define Doob’s (or McDiarmid’s) martingale
$d_{i}$ $=\mathbb{E}_{i}(X_{i}+\cdots+X_{n})-\mathbb{E}_{i-1}(X_{i}+\cdots+X_{n})$ (2.4)
@@
$S_{n}-\mathbb{E}S_{n}=\sum_{i=1}^{n}d_{i}\,\,,\quad E_{i-1}(d_{i})=0$
@@
$d_{i}$ $=X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)$ (2.5)
Given $u_{i-1}$, $E_{i}(X_{l})$ is a function of $X_{i}$. $E_{i-1}(X_{l})=E_{X_{i}}\{E_{i}(X_{l})\}$ where $E_{X_{i}}(\cdot)$ takes expectation with respect to $X_{i}$. We have
@@
Note, $\mathbb{E}_{i}(X_{\ell})\equiv\mathbb{E}(X_{\ell}\mid X_{i},u_{i-1})$. Given $u_{i-1}$, this is a function of $X_{i}$.
@@
With this notation, we have
@@
  • Poisson, Binomial, and Geometric distributions, for example, all satisfy the finite exponential moment condition, as does the Gaussian distribution.
@@
By definition (2.5),
$\mathbb{E}_{i-1}\left[e^{\delta|d_{i}|}\right]=\mathbb{E}_{i-1}\exp\left(\delta\,\Big|X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)\,\Big|\right)$
$\leq\mathbb{E}_{i-1}\Big[\exp\Big(\delta|X_{i}|+\delta|\mathbb{E}_{i-1}(X_{i})|+\delta\sum_{\ell=i+1}^{n}\big|\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big|\Big)\Big]$
@@
$\leq e^{\mathbb{E}_{i-1}(\delta|X_{i}|)}\,\mathbb{E}_{i-1}\Big[\exp\Big(\delta|X_{i}|+\delta|X_{i}|L\frac{M}{N}+\delta\,\mathbb{E}_{i-1}(\,|X_{i}|\,)\cdot L\frac{M}{N}\Big)\Big]$
$=\exp\left(\mathbb{E}_{i-1}\delta|X_{i}|\,+\,\delta L\frac{M}{N}\mathbb{E}_{i-1}|X_{i}|\right)\cdot\mathbb{E}_{i-1}\left[\exp\left(\delta|X_{i}|\Big(1+L\frac{M}{N}\Big)\right)\right]$
$=\exp\big(\mathbb{E}_{i-1}\delta\cdot A|X_{i}|\big)\cdot\mathbb{E}_{i-1}(e^{\delta A|X_{i}|})\leq K^{2}~~~~\mbox{(Jensen's inequality and the hypothesis)}$
Applying the exponential Markov inequality $\Pr[X\geq\lambda]\leq e^{-t\lambda}E(e^{tX})$ and noticing
$\mathbb{E}_{i-1}(d_{i})=0$
it follows
$\Pr\big(S_{n}-\mathbb{E}S_{n}\geq\lambda\,\big)$
$\leq e^{-t\lambda}\,\mathbb{E}\!\left[e^{\,t(S_{n}-\mathbb{E}S_{n})}\right]=e^{-t\lambda}\,\mathbb{E}\!\left[e^{t(d_{1}+d_{2}+\cdots+d_{n})}\right]=e^{-t\lambda}\,\mathbb{E}\!\left[e^{t(d_{1}+\cdots+d_{n-1})}\;\mathbb{E}_{n-1}\left(e^{td_{n}}\right)\right]$
[…]
$\leq e^{-t\lambda}\,e^{\,t^{2}K^{2}/\delta^{2}}\;\mathbb{E}\!\left[e^{t(d_{1}+\cdots+d_{n-2})}\;\mathbb{E}_{n-2}\!\left(e^{td_{n-1}}\right)\right]$
@@
  • The condition of Lipschitz continuity (2.10) may be interpreted as “continuity” of the trajectory with respect to the history; i.e. small differences in initial position do not lead to vehement differences in the trajectory.
@@
    and therefore $d_{i}$ (see, (2.5)) is reduced to $X_{i}-\mathbb{E}X_{i}$ (independence) and to $X_{i}$ (martingale difference), respectively. In these cases, the condition of bounded difference or the existence of exponential moment alone suffices to imply the concentration inequalities.
@@
The following theorem provides easily verifiable conditions for concentration inequality in general cases where
@@
i.e. $\frac{\partial y}{\partial t}$ is not constant with $u_{i}$ varying.
@@
Let $\mu_{i}^{\prime}$ denote $\frac{d\mu_{i}(u_{i})}{du_{i}}$, $\nu_{i}^{\prime}$ denote $\frac{d\nu_{i}(u_{i})}{du_{i}}$, $\mu_{i}^{\prime\prime}$ the second order derivative, and so on. We shall use term “smooth function” to refer that (a) the first several order derivatives (with respective to $u_{i}$, briefly $u$) exist and (b) they are uniformly bounded (in other words, the domain in question lies sufficiently away from singularities).
@@
In the notation of partial derivatives, the condition in Theorem 2.2 states that
@@
are uniformly bounded in the domain of smoothness (i.e., away from singularities), in addition to the exponential moment condition. These conditions are practically easy to verify because we can examine $\bar{u}$, the mean of $u$, when in particular it satisfies a partial differential equation (PDE). From the PDE we have every partial derivatives of $\bar{u}$, and $u$ is close to $\bar{u}$ with exponentially rare and negligible exceptions. Note, the condition (2.10) says that small changes in distant past generations won’t lead to violently change in the current step size. Theorem 2.2 tells us, in terms of PDE of $\bar{u}$, what local behavior of $u$ can meet (2.10), for the concentration inequality to hold.
@@
We only need to show Lipschitz continuity (2.10) holds, given the smoothness of $\mu_{i}$ and $\nu_{i}$. [^3] In most cases of interest, when $\mathbb{E}(X|u)$ is smooth, $\mathbb{E}(X^{2}|u)$ is smooth as well. Intuitively, the requirement that $\mathbb{E}(X^{2}|u)$ be smooth is the price we pay for not assuming boundedness of $X$. The proof is proceeded by induction. The base case holds, i.e. $|\mathbb{E}(X_{i+1}|u_{i})-\mathbb{E}(X_{i+1}|u_{i}^{\prime})|\leq L|u_{i}-u_{i}^{\prime}|$, because of the smoothness of $\mu_{i}$.
@@
$\mathbb{E}(X_{n+1})=\mu_{n}(\bar{u}_{n})~+\frac{1}{2}\mu_{n}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}+\frac{C}{N^{3/2}}$ (2.12)
@@
$\mathbb{E}(X_{n+1}^{2})=$ $\nu_{n}(\bar{u}_{n})+\frac{1}{2}\nu_{n}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}+~\frac{C}{N^{3/2}}$ (2.13)
@@
the following recurrence equation about $\mathrm{Var}_{n+1}$ holds,
@@
We first estimate the second term in the right hand-side of (2.14).
$\mathbb{E}\left[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})\right]=$ $\frac{1}{N}\mathbb{E}[X_{n+1}(u_{n}-\bar{u}_{n})]-\bar{x}_{n+1}(\bar{u}_{n}-\bar{u}_{n})$
@@
Again applying Taylor expansion to $\mu_{n}(u_{n})$, i.e.
@@
we have, with the term of $\sim\mathbb{E}(u_{n}-\bar{u}_{n})^{4}$ neglected,
@@
where the term of $\mathbb{E}(u_{n}-\bar{u}_{n})^{3}$ is neglected.
@@
$\mathbb{E}(X_{n+1}^{2})-\overline{X}_{n+1}^{2}=$ $\nu_{{n}}(\bar{u}_{n})+\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}-[\mu_{{n}}(\bar{u}_{n})+\frac{1}{2}\mu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}]^{2}$
@@
By substituting $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$ and $\mathbb{E}[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})]$ into (2.14), we have a recurrence relation for $n\geq i$,
@@
and $\mathrm{Var}_{i}=0$. With the argument $\bar{u}_{n}$” omitted for brevity:
@@
Then, the above recursion for $\mathrm{Var}$ gives
$\mathrm{Var}_{n+1}=$ $\sum_{j=i+1}^{n-1}b_{j}\cdot a_{j+1}...a_{n}+b_{n}=\sum_{j=i+1}^{n-1}b_{j}\cdot a_{j+1}...a_{n}+\frac{1}{N^{2}}(\nu_{n}-\mu_{n}^{2})$
$=$ $\frac{1}{N^{2}}\left[\sum_{j=i+1}^{n-1}(\nu_{j}-\mu_{j}^{2})\prod_{k=j+1}^{n}\left(1+\frac{1}{N}\Big(2\mu_{k}^{\prime}+\frac{1}{2N}\nu_{k}^{\prime\prime}-\frac{1}{N}\mu_{k}\mu_{k}^{\prime\prime}\Big)\right)+\nu_{n}-\mu_{n}^{2}\right]$
$=$ $\frac{1}{N^{2}}\left[\sum_{j=i+1}^{n-1}\left(\nu_{j}-\mu_{j}^{2}\right)\prod_{k=j+1}^{n}\left(1+\frac{2\mu_{k}^{\prime}}{N}\right)\right]$
(higher order terms are neglected)
Given above, it is easy to check that, in the domain where $\mu_{j}^{\prime}$, $\mu_{j}^{\prime\prime}$ and $\nu_{j}^{\prime\prime}$ are bounded, j=1,2nj=1,2...n, $\mathrm{Var}_{n+1}$ is bounded by a constant of $O(1/N)$. Further, by induction, $\frac{d\mathrm{Var}_{n+1}}{du}$ exists and its order is $O(1/N)$.
From the base that $\Delta u_{i}=u_{i}-u_{i}^{\prime}$, and the inductive hypothesis that $\Delta\bar{u}_{n-1}$ is small, it follows from (2.12) that
$\Delta\mathbb{E}(X_{n})$ $=\Delta\mu_{n-1}~+~\Delta\frac{1}{2}\mu_{n-1}^{\prime\prime}\mathrm{Var}_{n-1}~+~\frac{C}{N^{3/2}}$
@@
where the 3rd term in (2.12), of order $N^{-3/2}$, is omitted since it yields $\sim\displaystyle O(N^{-3/2})\Delta\bar{u}_{n-1}$ in the above equation. From $u_{n}=u_{n-1}+X_{n}/N$ and then
$\bar{u}_{n}=\bar{u}_{n-1}+\frac{1}{N}E(X_{n})$
we have
$\Delta\bar{u}_{n}=\Delta\bar{u}_{n-1}+\frac{1}{N}\Delta E(X_{n})=\Delta\bar{u}_{n-1}+\frac{1}{N}\Delta\bar{u}_{n-1}\Big(\mu_{n-1}^{\prime}+O\left(N^{-1}\right)+O(\Delta\bar{u}_{n-1})\Big)$
@@
$\Delta\bar{u}_{n}=\Delta\bar{u}_{n-1}\left(1+\frac{1}{N}\mu_{n-1}^{\prime}+o\Big(\frac{1}{N}\Big)\right)$
Iterating on $n$ yields
$|\,\mathbb{E}(X_{n}\mid u_{i})-\mathbb{E}(X_{n}\mid u_{i}^{\prime})\,|\leq L|u_{i}-u_{i}^{\prime}|\quad\text{(where $u_{i}-u_{i}^{\prime}=\Delta u_{i}$)}$
This concludes the proof. ∎
We see from the above proof that Lipschitz continuity of $\mathbb{E}(f(u_{n}))$ is ultimately down to the existence of the first several order derivatives of $f$.

### v6

@@ 1.1 Concentration inequalities
For a long time, there had been no substantive improvement in the eliminating boundedness condition (the trivial idea of truncation gives weak results, though) [^2] In fact, it was not until 1995 that the first improvement in Hoeffding inequality was made [42]. The reason might be two folds. First, there lacked motivation. Second, nifty techniques used in the pioneer work and the later literatures as well to achieve finer results somehow concealed the essential reason why the inequalities hold.. In early 2000s [39] relaxed the boundedness to finite exponential moments; i.e. for a $\delta>0,\mathbb{E}(e^{\delta X_{i}})$ is uniformly bounded for all $i$, and gave $\Pr(|S_{n}-\mathbb{E}(S_{n})|\geq\lambda|)\leq e^{-c\big(\frac{\lambda^{2}}{n}\big)^{1/3}}$— not Chernoff bound. Later, motivated by the study of the free energy of a directed polymer in a random environment, [34] gave Chernoff bound under the hypothesis that, for a $\delta>0,\mathbb{E}(e^{\delta X_{i}}\mid\mathcal{F}_{i-1})$ is bounded for all $i$ (where $\mathcal{F}_{i-1}$ is a filtration for the martingale difference sequence $X_{i}$). Other improvements of concentration inequalities in relaxing boundedness condition can be found in [8] [20] [26] [21]. However, to our knowledge, no existing work eliminates both the martingale and boundedness conditions while achieving sub-Gaussian concentration; related work by Wormald (1995) is discussed later. The first major contribution of this paper is to do so; informally, under the condition of “finite” increment, the small per-step effect in a random process can be controlled by the Lipschitz continuity of the mean increment with respect to the history, not necessarily by “almost independence”.
@@ 1.2 Notation and contributions
$\frac{\partial y}{\partial t}:=\frac{\mathbb{E}_{i}(y_{t})-\mathbb{E}_{i}(y_{t-1})}{1/N}=\mathbb{E}_{i}(Y_{t})-\mathbb{E}_{i}(Y_{t-1})=E_{i}(X_{t})\;(\equiv\mathbb{E}(X_{t}\mid u_{i}))$ (1.2)
@@
While classical concentration inequalities require the condition (1.4), one of the primary contributions of this work is the relaxation of this classical constraint to the condition:
@@ 2 Extension of Hoeffding-Azuma inequality
$\exp({\sim nt^{2}-t\lambda}),\quad$
for $\lambda\sim n^{(1+p)/2}$ with $0<p<1$, which can be made exponentially small—of order $\exp(-\Theta(n^{p}))$—by choosing sufficiently small $t$. Intuitively, provided $\lambda$ is not too small, the exponent can be made exponentially small by choosing small $t$, as $t^{2}$ is of higher order than $t$.
@@
The emphasis in this exposition is to show that the condition of $X_{i}<\infty$ can be relaxed to $E(e^{|X_{i}|})<\infty$ without compromising strength of results; from the practical point of view, the difference between the constant factors in the exponent, i.e., $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}$ versus $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$, is immaterial – both are square-exponential decay. In addition, it is shown that sub-Gaussian tail bounds can be obtained without “nifty” techniques.
Next, we identify weaker conditions that ensure concentration inequalities, beyond independence, martingale differences, and boundedness. To this end, define Doob’s (or McDiarmid’s) martingale,
$\Delta_{i}$ $=\mathbb{E}_{i}(X_{i}+\cdots+X_{n})-\mathbb{E}_{i-1}(X_{i}+\cdots+X_{n})$
@@
$S_{n}-\mathbb{E}S_{n}=\sum_{i=1}^{n}\Delta_{i}\,\,,\quad E_{i-1}(\Delta_{i})=0$
@@
$\Delta_{i}$ $=X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)$ (2.4)
Given $u_{i-1}$, $E_{i}(X_{l})$ is a function of $X_{i}$. $E_{i-1}(X_{l})=E_{X_{i}}\Big[E_{i}(X_{l})\Big]$ where $E_{X_{i}}(\cdot)$ takes expectation with respect to $X_{i}$. We have
@@
Note, $\mathbb{E}_{i}(X_{\ell})\equiv\mathbb{E}(X_{\ell}\mid X_{i},u_{i-1})$, which is a function of $X_{i}$ given $u_{i-1}$.
@@
With this notation and the tower property of conditional expectation, we have,
@@
  • Condition (2.7) in the Lemma 2.3 is equivalent to the following Lipschitz continuity
@@
    Informally, the condition (2.7) asserts that the drift is Lipschitz continuous with respect to history, including the initial state $S_{0}$. We will later show that it is true if the drift is smooth.
@@
  • Poisson(d)(d), Binomial(N,d/N)(N,d/N), and the variable occurrence distribution in SAT/COL models, for example, all satisfy the finite exponential moment condition, as does the Gaussian distribution.
@@
By definition (2.4), we have
$\mathbb{E}_{i-1}\left[e^{\delta|\Delta_{i}|}\right]=\mathbb{E}_{i-1}\exp\left(\delta\,\Big|X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)\,\Big|\right)$
$\leq\mathbb{E}_{i-1}\exp\left(\delta|X_{i}|+\delta|\mathbb{E}_{i-1}(X_{i})|+\delta\sum_{\ell=i+1}^{n}\left|\,\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\,\right|\right)$
$\leq e^{\mathbb{E}_{i-1}(\delta|X_{i}|)}\,\mathbb{E}_{i-1}\exp\left(\delta|X_{i}|+\delta\sum_{\ell=i+1}^{n}L\frac{|X_{i}|}{N}+\delta\sum_{\ell=i+1}^{n}L\frac{\mathbb{E}_{i-1}|X_{i}|}{N}\right)$
(Lemma 2.2)
$\leq e^{\mathbb{E}_{i-1}(\delta|X_{i}|)}\,\mathbb{E}_{i-1}\exp\left(\delta|X_{i}|+\delta|X_{i}|L\frac{M}{N}+\delta\,\mathbb{E}_{i-1}(\,|X_{i}|\,)\cdot L\frac{M}{N}\right)$
$=\exp\left(\delta\mathbb{E}_{i-1}|X_{i}|\,+\,\delta L\frac{M}{N}\mathbb{E}_{i-1}|X_{i}|\right)\cdot\mathbb{E}_{i-1}\exp\left(\delta|X_{i}|\Big(1+L\frac{M}{N}\Big)\right)$
$=\exp\big(\delta A\mathbb{E}_{i-1}|X_{i}|\big)\cdot\mathbb{E}_{i-1}\left[e^{\delta A|X_{i}|}\right]\leq K^{2}$
(Jensen’s inequality and the hypothesis)
Applying the exponential Markov inequality $\Pr[X\geq\lambda]\leq e^{-t\lambda}E(e^{tX})$ and noting
$\mathbb{E}_{i-1}(\Delta_{i})=0$
we obtain
$\Pr\big(S_{n}-\mathbb{E}S_{n}\geq\lambda\,\big)$ $\leq e^{-t\lambda}\,\mathbb{E}\!\left[e^{\,t(S_{n}-\mathbb{E}S_{n})}\right]$
$=e^{-t\lambda}\,\mathbb{E}\!\left[e^{t(\Delta_{1}+\Delta_{2}+\cdots+\Delta_{n})}\right]$
$=e^{-t\lambda}\,\mathbb{E}\!\left[e^{t(\Delta_{1}+\cdots+\Delta_{n-1})}\;\mathbb{E}_{n-1}\left(e^{t\Delta_{n}}\right)\right]$
@@
$\leq e^{-t\lambda}\,e^{\,t^{2}K^{2}/\delta^{2}}\;\mathbb{E}\!\left[e^{t(\Delta_{1}+\cdots+\Delta_{n-2})}\;\mathbb{E}_{n-2}\!\left(e^{t\Delta_{n-1}}\right)\right]$
@@
  • The condition of Lipschitz continuity (2.9) may be interpreted as “continuity” of the trajectory with respect to the history; i.e. small differences in initial position do not lead to drastic differences in the trajectory.
@@
    and therefore $\Delta_{i}$ (see, (2.4)) reduces to $X_{i}-\mathbb{E}X_{i}$ (independence) and to $X_{i}$ (martingale difference), respectively. In these cases, the condition of bounded difference or the existence of exponential moment alone suffices to imply the concentration inequalities.
@@
The following theorem provides readily verifiable conditions for concentration inequality in general settings where
@@
i.e. $\frac{\partial y}{\partial t}$ is not necessarily constant with $u_{i}$ varying.
@@
Let $\mu_{i}^{\prime}$ denote $\frac{d\mu_{i}(u_{i})}{du_{i}}$, $\nu_{i}^{\prime}$ denote $\frac{d\nu_{i}(u_{i})}{du_{i}}$, $\mu_{i}^{\prime\prime}$ the second order derivative, and so on. We shall use term “smooth function” to refer that (a) the first several order derivatives (with respective to $u_{i}$, briefly $u$) exist and (b) they are uniformly bounded (in other words, the domain in question is bounded away from the singularities).
@@
In terms of partial derivatives, the condition in Theorem 2.2 states that
@@
are uniformly bounded on the domain (i.e., away from singularities), in addition to the exponential moment condition. These conditions are practically easy to verify because we can examine $\bar{u}$, the mean of $u$, when in particular it satisfies a partial differential equation (PDE). From the PDE we have every partial derivatives of $\bar{u}$, and $u$ is close to $\bar{u}$ with exponentially rare and negligible exceptions. Note, the condition (2.9) says that small changes in distant past generations won’t lead to violently change in the current step size. Theorem 2.2 tells us, in terms of PDE of $\bar{u}$, what local behavior of $u$ can meet (2.9), for the concentration inequality to hold.
@@
We only need to show Lipschitz continuity (2.9) holds, given the smoothness of $\mu_{i}$ and $\nu_{i}$$\dagger$$\dagger\daggerIn$ most cases of interest, when $\mathbb{E}(X|u)$ is smooth, $\mathbb{E}(X^{2}|u)$ is smooth as well. Intuitively, regularity of $\mathbb{E}(X^{2}|u)$ is the price paid for dispensing with a boundedness assumption on $X$.
The proof is proceeded by induction. The base case holds, i.e. $|\mathbb{E}(X_{i+1}|u_{i})-\mathbb{E}(X_{i+1}|u_{i}^{\prime})|\leq L|u_{i}-u_{i}^{\prime}|$, because of the smoothness of $\mu_{i}$.
@@
$\mathbb{E}(X_{n+1})=\mu_{n}(\bar{u}_{n})~+\frac{1}{2}\mu_{n}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}+\sim\frac{1}{N^{3/2}}$ (2.11)
@@
$\mathbb{E}(X_{n+1}^{2})=$ $\nu_{n}(\bar{u}_{n})+\frac{1}{2}\nu_{n}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}+\sim\frac{1}{N^{3/2}}$ (2.12)
@@
the following recurrence equation about $\mathrm{Var}_{n}$ holds,
@@
Applying the law of total expectation with $\mathbb{E}[X_{n+1}\,|\,u_{n}]=\mu_{n}(u_{n})$, the cross term in the right hand-side of (2.13) is
$\mathbb{E}\left[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})\right]=$ $\frac{1}{N}\mathbb{E}\left[\,X_{n+1}(u_{n}-\bar{u}_{n})\,\right]-\mathbb{E}\left[\,\bar{x}_{n+1}(u_{n}-\bar{u}_{n})\,\right]$
@@
Now from the Taylor expansion of $\mu_{n}(u_{n})$ around $\bar{u}$, i.e.
@@
neglecting higher-order terms, $\mathbb{E}(u_{n}-\bar{u}_{n})^{3}$ and terms of order $\mathbb{E}(u_{n}-\bar{u}_{n})^{4}$, we have,
@@
where we have used $\mathbb{E}(u_{n}-\bar{u}_{n})=0$ and omitted the cubic terms as negligible.
@@
$\mathbb{E}(X_{n+1}^{2})-\overline{X}_{n+1}^{2}=$ $\nu_{{n}}(\bar{u}_{n})+\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}-\left[\mu_{{n}}(\bar{u}_{n})+\frac{1}{2}\mu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}\right]^{2}$
@@
For final variance recurrence, substituting $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$ and $\mathbb{E}[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})]$ back into (2.13), we have the recurrence relation for $n\geq i$,
@@
and the base case is $\mathrm{Var}_{i}=0$. With the argument $\bar{u}_{n}$” omitted for brevity:
@@
Then, the variance recurrence relation for $n\geq i$ takes the compact form
$\mathrm{Var}_{n+1}=a_{n}\cdot\mathrm{Var}_{n}+b_{n}$
It follows, for $n\geq i+1$
$\mathrm{Var}_{n}=$ $\mathrm{Var}_{i}\prod_{k=i}^{n-1}a_{k}+\sum_{j=i}^{n-1}\Bigg(b_{j}\prod_{k=j+1}^{n-1}a_{k}\Bigg)$
$=$ $\frac{1}{N^{2}}\left[\sum_{j=i}^{n-1}(\nu_{j}-\mu_{j}^{2})\prod_{k=j+1}^{n-1}\left(1+\frac{1}{N}\Big(2\mu_{k}^{\prime}+\frac{1}{2N}\nu_{k}^{\prime\prime}-\frac{1}{N}\mu_{k}\mu_{k}^{\prime\prime}\Big)\right)\right]$
$=$ $\frac{1}{N^{2}}\left[\sum_{j=i}^{n-1}\left(\nu_{j}-\mu_{j}^{2}\right)\prod_{k=j+1}^{n-1}\left(1+\frac{2\mu_{k}^{\prime}}{N}\right)\right]$
where higher order terms in $N^{-1}$ have been neglected.
[…]
From the inductive hypothesis that $|\Delta\bar{u}_{n-1}|\leq L_{n-1}|u_{i}-u_{i}^{\prime}|=O(\Delta u_{i})$, so that $\Delta\bar{u}_{n-1}$ is small, it follows from (2.11) that, at step $n$,
$\Delta\mathbb{E}(X_{n})=\Delta\mu_{n-1}~+~\Delta\frac{1}{2}\mu_{n-1}^{\prime\prime}\mathrm{Var}_{n-1}~+~\frac{C}{N^{3/2}}\,.$
Expanding,
@@
Here the third term in (2.11), of order $N^{-3/2}$, has been omitted since it contributes only $O(N^{-3/2})\Delta\bar{u}_{n-1}$ in the above expansion. From
$u_{n}=u_{n-1}+\frac{X_{n}}{N},\qquad\bar{u}_{n}=\bar{u}_{n-1}+\frac{1}{N}\mathbb{E}(X_{n})$
it follows
$\Delta\bar{u}_{n}$ $=\Delta\bar{u}_{n-1}+\frac{1}{N}\Delta\mathbb{E}(X_{n})$
$=\Delta\bar{u}_{n-1}+\frac{1}{N}\Delta\bar{u}_{n-1}\Big(\mu_{n-1}^{\prime}+O(N^{-1})+O(\Delta\bar{u}_{n-1})\Big)$
@@
$\Delta\bar{u}_{n}=\Delta\bar{u}_{n-1}\left(1+\frac{1}{N}\mu_{n-1}^{\prime}+\frac{o(1)}{N}\right)$
Iterating in $n$ yields
$\left|\mathbb{E}(X_{n}\mid u_{i})-\mathbb{E}(X_{n}\mid u_{i}^{\prime})\right|\leq L_{n}\left|u_{i}-u_{i}^{\prime}\right|,\quad u_{i}-u_{i}^{\prime}=\Delta u_{i},$
with $L_{n}=\Theta(L)$. This concludes the proof. ∎
From the above proof, we see that Lipschitz continuity of $\mathbb{E}(f(u))$ with respect to the process history is down to the existence of the first few derivatives of $f$ and of the variance of $u$.